OpenStudy (anonymous):

Are there any inflection points for f(x)=x^2/((x^2)-4)?

OpenStudy (anonymous):

derivate it twice and set the it equal to zero...in other words solve this equation\[f''(x)=0\]

OpenStudy (anonymous):

I got 32+24x^2 as the numerator

OpenStudy (anonymous):

let me check it

OpenStudy (anonymous):

quite right :) and that results in ????

OpenStudy (anonymous):

nothing?

OpenStudy (anonymous):

is there any solution for\[32+24x^2=0\]???

OpenStudy (anonymous):

I don't think so? o.O

OpenStudy (anonymous):

it gives \[x^2=-\frac{4}{3}\] which is not a true statement for real numbers

OpenStudy (anonymous):

So no inflection?

OpenStudy (anonymous):

yes...no inflection point

OpenStudy (anonymous):

hm...ok thanks :)

OpenStudy (anonymous):

yw